OpenStudy (anonymous):

stacy states that root39 + root9 - root 4 =root12 explain why tracy is wrong

OpenStudy (micahm):

this site might help http://www.freemathhelp.com/

OpenStudy (anonymous):

micahm if you cant help then leave the premises

OpenStudy (micahm):

this is helping

OpenStudy (vishweshshrimali5):

Hi Daftkillz

OpenStudy (perl):

for starters, it is not true generally that sqrt( a ) + sqrt(b) = sqrt( a + b )

OpenStudy (vishweshshrimali5):

Your question is: Why this statement is wrong: \[\large{\sqrt{39}+\sqrt{9}-\sqrt{4} = \sqrt{12}}\] Now, as @perl mentioned, \[\large{\sqrt{a}+\sqrt{b} \ne \sqrt{a+b}}\]

OpenStudy (confluxepic):

Are Stacey and Tracey the same person.

OpenStudy (confluxepic):

http://prntscr.com/51lfrk